Let us consider a country of 100 people. Each person can produce 5 units of Good X or 10 units of Good Y a. b. C. d, e. Draw the PPF of the country, with Good X on x-axis and Good Y on y-axis. Write out the function of the PPF. What is the slope of the PPF? What does the slope of the PPF mean? if the market prices are PX-Py-1, at which point will the economy produce? (hard) Suppose another group of 100 people now immigrate to this country (so the country now has 200 people in total). In this group, each person can produce 10 units of Good X or 10 units of Good Y. Re-draw the PPF of the country. And write out the function of the PPF.
Answer #1

Answer

a)

b)

If 100 people were used on producing X then country can produce 100*5 = 500 and If If 100 people were used on producing Y then country can produce 100*10 = 1000

Hence we have 2 points (0 , 1000) and (500 , 0)

Hence as Opportunity cost of producing Y is constant in terms of X. This implies PPF is a straight Line.

Equation is given by:

(Y - 0) = ((0 - 1000)/(500 - 0))(X - 500)

=> 2X + Y = 1000 ------------------------Equation of PPF

c)

Slope of PPF = dY/dX = -2

Slope of PPF means how much Y country have to sacrifice in order to produce 1 more unit of X. It is also called Marginal Rate of Transformation

d)

Market Prices of X and Y are same. Hence GDP will maximized If they produced only Y. Hence country will produce at Y intercept i.e. Point P(see above graph)